### v1

# On Schultz’s generalization of Borweins’ cubic identity

Abstract.
Around 1991, J.M. and P.B. Borwein established a cubic analogue of Jacobi’s fundamental identity for theta functions. Their identity serves as the foundation for the subsequent development by B.C. Berndt, S. Bhargava, and F.G. Garvan of Ramanujan’s theory of elliptic functions to the cubic base. In 2013, D. Schultz discovered an identity for theta series in three variables which generalizes the Borweins’ identity. In this article, we revisit Schultz’s identity and present two distinct approaches to its derivation. Our investigation not only provides new proofs but also yields several identities of a similar type. Furthermore, this study enables us to construct new two-variable generalizations of Jacobi’s original classical identity.
2020 Mathematics Subject Classification:
33E05, 11F27
The second author is partially supported by the Ministry of Education, Singapore, Academic Research Fund, Tier 1 (RG15/23).
The third author is partially supported by the National Natural Science Foundation of China (Grant No. 12371328) and the Ministry of Education, Singapore, Academic Research Fund, Tier 1 (RG15/23).
1. Introduction
Around 1991, J.M. Borwein and P.B. Borwein [4, (2.3)] published the identity
$\left(\sum_{m,n\in\mathbf{Z}}\omega^{m-n}q^{m^{2}+mn+n^{2}}\right)^{3}$ $+\left(\sum_{m,n\in\mathbf{Z}}q^{(m+1/3)^{2}+(m+1/3)(n+1/3)+(n+1/3)^{2}}\right)^{3}$
$\qquad\qquad=\left(\sum_{m,n\in\mathbf{Z}}q^{m^{2}+mn+n^{2}}\right)^{3},$ (1.1)
where $\omega=e^{2\pi i/3}$. Their identity is a cubic analogue of Jacobi’s classical identity [14, p.90, (14)], [13, Theorem 3.7, Corollary 3.8] given by
$\left(\sum_{m,n\in\mathbf{Z}}(-1)^{m+n}q^{m^{2}+n^{2}}\right)^{2}+\left(\sum_{m,n\in\mathbf{Z}}q^{(m+1/2)^{2}+(n+1/2)^{2}}\right)^{2}=\left(\sum_{m,n\in\mathbf{Z}}q^{m^{2}+n^{2}}\right)^{2}.$ (1.2)
The Borweins used (1.1) to show that
$\sum_{m,n\in\mathbf{Z}}q^{m^{2}+mn+n^{2}}=\,_{2}F_{1}\left(\dfrac{1}{3},\dfrac{2}{3};1;\kappa(q)\right)$ (1.3)
where
$\kappa(q)=\left(\dfrac{\displaystyle\sum_{m,n\in\mathbf{Z}}q^{(m+1/3)^{2}+(m+1/3)(n+1/3)+(n+1/3)^{2}}}{\displaystyle\sum_{m,n\in\mathbf{Z}}q^{m^{2}+mn+n^{2}}}\right)^{3},$
$_{2}F_{1}\left(a,b;c;z\right)=\sum_{j=0}^{\infty}\dfrac{(a)_{j}(b)_{j}}{(c)_{j}}\dfrac{z^{j}}{j!}$
and
$(a)_{j}=\begin{cases}\qquad 1\quad&\text{when $j=0$,}\\ \displaystyle\prod_{k=1}^{j}(a+k-1)\quad&\text{when $j\in\mathbf{Z}^{+}$.}\end{cases}$
Borweins’ inversion formula (1.3) is crucial in proving Ramanujan’s identity [17, (31)], [5]
$\dfrac{27}{4\pi}=\sum_{j=0}^{\infty}(15j+2)\dfrac{(1/3)_{j}(2/3)_{j}(1/2)_{j}}{(1)_{j}^{3}}\left(\dfrac{2}{27}\right)^{j}.$
Using Borweins’ identities (1.1) and (1.3), B.C. Berndt, S. Bhargava and F.G. Garvan [1] developed Ramanujan’s theory of elliptic functions to the cubic base. An alternative approach to Ramanujan’s theory of elliptic functions to the cubic base without the use of (1.1) was recently discovered by H.H. Chan and Z.G. Liu [10].
In 2013, D. Schultz [19] published the identity given by
$\left(\sum_{m,n\in\mathbf{Z}}\omega^{m-n}q^{(m+1/3)^{2}+(m+1/3)(n+1/3)+(n+1/3)^{2}}x^{m+1/3}y^{n+1/3}\right)^{3}$
$+\left(\sum_{m,n\in\mathbf{Z}}q^{m^{2}+mn+n^{2}}x^{m}y^{n}\right)^{3}=\left(\sum_{m,n\in\mathbf{Z}}\omega^{m-n}q^{m^{2}+mn+n^{2}}x^{m}y^{n}\right)^{3}$
$\qquad\qquad+\left(\sum_{m,n\in\mathbf{Z}}q^{(m+1/3)^{2}+(m+1/3)(n+1/3)+(n+1/3)^{2}}x^{m+1/3}y^{n+1/3}\right)^{3}.$ (1.4)
Schultz’s identity turns out to be a two-variable generalization of the (1.1). To see this, we need to show that
$\sum_{m,n\in\mathbf{Z}}\omega^{m-n}q^{(m+1/3)^{2}+(m+1/3)(n+1/3)+(n+1/3)^{2}}=0.$ (1.5)
Identity (1.5) is equivalent to
$S=\sum_{m,n\in\mathbf{Z}}\omega^{m-n}q^{m^{2}+mn+n^{2}+m+n}=0.$
Replacing $(m,n)$ by (n,mn1)(n,-m-n-1), we deduce that
$S=\omega^{-1}.S$
This implies that $S=0$ and the proof of (1.5) is complete. For other proofs of (1.5), see [23, Corollary] and [20, Section 2.3.3]. Substituting $x=y=1$ in (1.4) and using (1.5), we arrive at (1.1).
The first proof of (1.4) was provided by Schultz [19, p. 622]. He also mentioned, without giving explicit details, four approaches [19, p. 663] to establish the identity. The only other proof of (1.4) which we are aware of is due to X.R. Ma and R.Z. Wei [16, Corollary 4.12].
In this article, we give two new proofs of (1.4). Our first approach is motivated by the work of X.M. Yang [24]. In Section 2, we show that (1.4) follows from one of her identities [24, (3.8)]. In Section 4, we use transformation formulas for theta functions and the identities in Sections 2 and 3 to derive analogues of (1.4).
Our second proof of (1.4) connects the identity to the study of Macdonald’s identities [9] and this is presented in Section 4.
Motivated by (1.4), we found an analogue of (1.4) for Jacobi’s identity (1.2) given by
$\left(\sum_{m,n\in\mathbf{Z}}(-1)^{m+n}q^{(m+1/2)^{2}+(n+1/2)^{2}}x^{m+1/2}y^{n+1/2}\right)^{2}+\left(\sum_{m,n\in\mathbf{Z}}q^{m^{2}+n^{2}}x^{m}y^{n}\right)^{2}$
$=\left(\sum_{m,n\in\mathbf{Z}}q^{(m+1/2)^{2}+(n+1/2)^{2}}x^{m+1/2}y^{n+1/2}\right)^{2}+\left(\sum_{m,n\in\mathbf{Z}}(-1)^{m+n}q^{m^{2}+n^{2}}x^{m}y^{n}\right)^{2}.$ (1.6)
Since
$\sum_{j=-\infty}^{\infty}(-1)^{j}q^{(j+1/2)^{2}}$ $=\sum_{\ell=-\infty}^{\infty}(-1)^{\ell-1}q^{(\ell-1/2)^{2}}$
$=\sum_{k=-\infty}^{\infty}(-1)^{-k-1}q^{(-k-1/2)^{2}}$
$=-\sum_{k=-\infty}^{\infty}(-1)^{k}q^{(k+1/2)^{2}},$
we conclude that
$\sum_{m,n\in\mathbf{Z}}(-1)^{m+n}q^{(m+1/2)^{2}+(n+1/2)^{2}}=0.$ (1.7)
By (1.7), we see that (1.2) follows from (1.6) with the substitution $x=y=1$.
In Section 5, we derive several extensions of (1.6). In Section 6, we derive some identities equivalent to (1.1). The identities presented in these last two sections are motivated by the recent discoveries of H.H. Chan, S.H. Chan and P. Sole [8].
2. Jacobi’s theta functions and (1.4)
Throughout this work, $q=e^{\pi i\tau}$ where the imaginary part of $\tau$ is positive. The Jacobi theta functions are defined by
$\vartheta_{1}(z|\tau)=-i\sum_{n=-\infty}^{\infty}(-1)^{n}q^{(2n+1)^{2}/4}e^{(2n+1)iz},$
$\vartheta_{2}(z|\tau)=\sum_{n=-\infty}^{\infty}q^{(2n+1)^{2}/4}e^{(2n+1)iz},$
$\vartheta_{3}(z|\tau)=\sum_{n=-\infty}^{\infty}q^{n^{2}}e^{2niz}$
and
$\vartheta_{4}(z|\tau)=\sum_{n=-\infty}^{\infty}(-1)^{n}q^{n^{2}}e^{2niz}.$
A straightforward calculation also yields the following expression for $\vartheta_{1}$:
$\vartheta_{1}(z|\tau)=2\sum_{n=0}^{\infty}(-1)^{n}q^{(2n+1)^{2}/4}\sin(2n+1)z.$ (2.1)
The infinite product representations of $\vartheta_{1}(z|\tau)$ and $\vartheta_{3}(z|\tau)$ ([22, p. 469], [7, p. 30]), which are equivalent to Jacobi’s triple product identity, are given by
$\vartheta_{1}(z|\tau)$ $=2q^{1/4}(\sin z)\prod_{n=1}^{\infty}(1-q^{2n})(1-q^{2n}e^{2iz})(1-q^{2n}e^{-2iz})$ (2.2)
and
$\vartheta_{3}(z|\tau)$ $=\prod_{n=1}^{\infty}(1-q^{2n})(1+q^{2n-1}e^{2iz})(1+q^{2n-1}e^{-2iz}).$
The Jacobi theta function $\vartheta_{1}(z|\tau)$ satisfies
$\vartheta_{1}(z|\tau)=-\vartheta_{1}(z+\pi|\tau)=-qe^{2iz}\vartheta_{1}(z+\pi\tau|\tau)$ (2.3)
and
$\vartheta_{1}(z+(\pi+\pi\tau)/2|\tau)=q^{-1/4}e^{-iz}\vartheta_{3}(z|\tau).$ (2.4)
Let
$\Lambda=\left\{x\pi+y\pi\tau\big|0\leq x<1,0\leq y<1\right\}$
be the fundamental period parallelogram. In $\Lambda$, $\vartheta_{1}(z|\tau)$ has a simple zero at $z=0$ and $\vartheta_{3}(z|\tau)$ has a simple zero at $z=(\pi+\pi\tau)/2$. Using (2.3) and (2.4), it is immediate that
$\vartheta_{1}(3z|3\tau),\quad e^{2iz}\vartheta_{1}(3z+\pi\tau|3\tau)\quad\text{and}\quad e^{-2iz}\vartheta_{1}(3z-\pi\tau|3\tau)$
satisfy the relations
$f(z)=-f(z+\pi)=-q^{3}e^{6iz}f(z+\pi\tau).$ (2.5)
The next lemma plays an important role in our approach to (1.4).
Lemma 2.1.
The functions $\vartheta_{1}(3z|3\tau)$, $e^{2iz}\vartheta_{1}(3z+\pi\tau|3\tau)$ and $e^{-2iz}\vartheta_{1}(3z-\pi\tau|3\tau)$ are linearly independent over $\mathbf{C}$.
Proof.
Suppose there are complex numbers $c_{1},c_{2}$ and $c_{3}$ such that
$c_{1}\vartheta_{1}(3z|3\tau)+c_{2}e^{2iz}\vartheta_{1}(3z+\pi\tau|3\tau)+c_{3}e^{-2iz}\vartheta_{1}(3z-\pi\tau|3\tau)=0.$ (2.6)
If we set $z=0$ in (2.6) and use the fact that $\vartheta_{1}(0|3\tau)=0$, we find that
$(c_{2}-c_{3})\vartheta_{1}(\pi\tau|3\tau)=0.$
Since $\vartheta_{1}(\pi\tau|3\tau)=iq^{-1/6}\eta(\tau)\not=0$, we deduce that $c_{2}=c_{3}$. Substituting $c_{2}=c_{3}$ into (2.6), we obtain
$c_{1}\vartheta_{1}(3z|3\tau)+c_{2}\left(e^{2iz}\vartheta_{1}(3z+\pi\tau|3\tau)+e^{-2iz}\vartheta_{1}(3z-\pi\tau|3\tau)\right)=0.$ (2.7)
Next, we let $z=\pi/3$ in (2.7). By (2.1), $\vartheta_{1}(\pi|3\tau)=0$ and therefore,
$i\sqrt{3}\vartheta_{1}(\pi\tau|3\tau)c_{2}=0,$
which implies that $c_{2}=0$ since $\vartheta_{1}(\pi\tau|3\tau)\neq 0$. Therefore, $c_{2}=c_{3}=0$. Identity (2.6) reduces to
$c_{1}\vartheta_{1}(3z|3\tau)=0$
which implies that $c_{1}=0$ since $\vartheta_{1}(3z|3\tau)\neq 0$. The three functions are therefore linearly independent over $\mathbf{C}$. ∎
Let
$T(z,x,y|\tau)=\vartheta_{1}(z+x|\tau)\vartheta_{1}(z+y|\tau)\vartheta_{1}(z-x-y|\tau).$
The function $T(z,x,y|\tau)$ is constructed so that it satisfies (2.5) since
$T(z,x,y|\tau)=-q^{3}e^{2i(z+x+z+y+z-x-y)}T(z+\pi\tau,x,y|\tau)=-q^{3}e^{6iz}T(z+\pi\tau,x,y|\tau).$
Therefore, the functions
$\vartheta_{1}(3z|3\tau)/T(z,x,y|\tau),$
$e^{2iz}\vartheta_{1}(3z+\pi\tau|3\tau)/T(z,x,y|\tau)$
and
$e^{-2iz}\vartheta_{1}(3z-\pi\tau|3\tau)/T(z,x,y|\tau)$
are all elliptic functions with periods $\pi$ and $\pi\tau$ and have 3 poles in $\Lambda$. Since an elliptic function with a simple pole must be a constant, we conclude that there exists $c,\lambda_{1},\lambda_{2}$ and $\lambda_{3}$ which are dependent on $x$ and $y$ but independent of $z$ such that
$cT(z,x,y|\tau)=\lambda_{1}\vartheta_{1}(3z|3\tau)+\lambda_{2}e^{2iz}\vartheta_{1}(3z+\pi\tau|3\tau)+\lambda_{3}e^{-2iz}\vartheta_{1}(3z-\pi\tau|3\tau).$
Note that by Lemma 2.1, $c\neq 0$. By replacing $\lambda_{j}$ by $\lambda_{j}/c$, we find that
$\vartheta_{1}$ $(z+x|\tau)\vartheta_{1}(z+y|\tau)\vartheta_{1}(z-x-y|\tau)$
$=\lambda_{1}\vartheta_{1}(3z|3\tau)+\lambda_{2}e^{2iz}\vartheta_{1}(3z+\pi\tau|3\tau)+\lambda_{3}e^{-2iz}\vartheta_{1}(3z-\pi\tau|3\tau).$ (2.8)
Replacing $z$ by $z+(\pi+\pi\tau)/2$ and using (2.4), we may rewrite (2.8) as
$\vartheta_{3}(z+x|\tau)\vartheta_{3}(z+y|\tau)\vartheta_{3}(z-x-y|\tau)$
$\qquad\qquad=-\lambda_{1}\vartheta_{3}(3z|3\tau)+\lambda_{2}e^{2iz}\vartheta_{3}(3z+\pi\tau|3\tau)+\lambda_{3}e^{-2iz}\vartheta_{3}(3z-\pi\tau|3\tau).$ (2.9)
Identity (2.8) written in the form (2.9) allows us to determine $\lambda_{1},\lambda_{2}$ and $\lambda_{3}$. Substituting the series representation of $\vartheta_{3}(z|\tau)$ into (2.9) and equating the constant terms, the coefficients of $e^{2iz}$ and $e^{-2iz}$, we deduce that
$\lambda_{1}$ $=-\sum_{\begin{subarray}{c}k,\ell,m\in\mathbf{Z}\\ k+\ell+m=0\end{subarray}}q^{\ell^{2}+k^{2}+m^{2}}e^{2i((\ell-m)x+(k-m)y)}$
$=-\sum_{\begin{subarray}{c}k,\ell\in\mathbf{Z}\end{subarray}}q^{2(\ell^{2}+\ell k+k^{2})}e^{2i(2\ell+k)x}e^{2i(2k+\ell)y},$
$\lambda_{2}$ $=\sum_{\begin{subarray}{c}k,\ell,m\in\mathbf{Z}\\ k+\ell+m=1\end{subarray}}q^{\ell^{2}+k^{2}+m^{2}}e^{2i((\ell-m)x+(k-m)y)}$
$=q^{1/3}e^{-2ix}e^{-2iy}\sum_{\begin{subarray}{c}k,\ell\in\mathbf{Z}\end{subarray}}q^{2(\ell^{2}+\ell k+k^{2}-k-\ell+1/3)}e^{2i(2\ell+k)x}e^{2i(2k+\ell)y}$
and
$\lambda_{3}$ $=\sum_{\begin{subarray}{c}k,\ell,m\in\mathbf{Z}\\ k+\ell+m=-1\end{subarray}}q^{\ell^{2}+k^{2}+m^{2}}e^{2i((\ell-m)x+(k-m)y)}$
$=q^{1/3}e^{2ix}e^{2iy}\sum_{\begin{subarray}{c}k,\ell\in\mathbf{Z}\end{subarray}}q^{2(\ell^{2}+\ell k+k^{2}+k+\ell+1/3)}e^{2i(2\ell+k)x}e^{2i(2k+\ell)y}.$
The identities we derived are recorded as follow:
Theorem 2.1.
Let $z,x,y\in\mathbf{C}$ and $q=e^{\pi i\tau}$ with $\text{Im}\,\tau>0$. Let
$A(x,y|\tau)=\sum_{\begin{subarray}{c}k,\ell\in\mathbf{Z}\end{subarray}}q^{2(\ell^{2}+\ell k+k^{2})}e^{2ix(2\ell+k)}e^{2iy(2k+\ell)}$ (2.10)
and
$C(x,y|\tau)=e^{2ix}e^{2iy}\sum_{\begin{subarray}{c}k,\ell\in\mathbf{Z}\end{subarray}}q^{2(\ell^{2}+\ell k+k^{2}+k+\ell+1/3)}e^{2ix(2\ell+k)}e^{2iy(2k+\ell)}.$ (2.11)
Then
$\vartheta_{1}(z+x|\tau)\vartheta_{1}(z+y|\tau)\vartheta_{1}(z-x-y|\tau)=-A(x,y|\tau)\vartheta_{1}(3z|3\tau)$ (2.12)
$\quad+q^{1/3}C(-x,-y|\tau)e^{2iz}\vartheta_{1}(3z+\pi\tau|3\tau)+q^{1/3}C(x,y|\tau)e^{-2iz}\vartheta_{1}(3z-\pi\tau|3\tau),$
and
$\vartheta_{3}(z+x|\tau)\vartheta_{3}(z+y|\tau)\vartheta_{3}(z-x-y|\tau)=A(x,y|\tau)\vartheta_{3}(3z|3\tau)$ (2.13)
$\quad+q^{1/3}C(-x,-y|\tau)e^{2iz}\vartheta_{3}(3z+\pi\tau|3\tau)+q^{1/3}C(x,y|\tau)e^{-2iz}\vartheta_{3}(3z-\pi\tau|3\tau).$
We are now ready to prove (1.4).
Proof of (1.4).
First, let $z=-\pi\tau/3$ in (2.12) to deduce that
$\vartheta_{1}(-\pi\tau/3+x|\tau)\vartheta_{1}(-\pi\tau/3+y|\tau)\vartheta_{1}(-\pi\tau/3-x-y|\tau)\$
$\qquad\qquad\qquad=-A(x,y|\tau)\vartheta_{1}(-\pi\tau|3\tau)+qC(x,y|\tau)\vartheta_{1}(-2\pi\tau|3\tau).$
By (2.2), we know that
$\vartheta_{1}(-\pi\tau|3\tau)=-iq^{-1/3}\eta(\tau)\quad\text{and}\quad\vartheta_{1}(-2\pi\tau|3\tau)=-iq^{-4/3}\eta(\tau),$ (2.14)
where
$\eta(\tau)=q^{1/12}\prod_{k=1}^{\infty}(1-q^{2k}).$
Therefore,
$\dfrac{iq^{1/3}}{\eta(\tau)}\vartheta_{1}(-\pi\tau/3+x|\tau)\vartheta_{1}(-\pi\tau/3+y|\tau)\vartheta_{1}(-\pi\tau/3-x-y|\tau)=A(x,y|\tau)-C(x,y|\tau),$
which, after replacing $\tau$ by $3\tau$, implies that
$A(x,y|3\tau)-C(x,y|3\tau)=-\dfrac{iq}{\eta(3\tau)}\vartheta_{1}(-\pi\tau+x|3\tau)\vartheta_{1}(-\pi\tau+y|3\tau)\vartheta_{1}(-\pi\tau-x-y|3\tau).$
Using (2.2), we deduce that
$A(x,y|3\tau)-C(x,y|3\tau)=\prod_{k=1}^{\infty}(1-q^{6k})^{2}(1-q^{6k-2}e^{2ix})(1-q^{6k-4}e^{-2ix})$
$\qquad\times\prod_{k=1}^{\infty}(1-q^{6k-2}e^{2iy})(1-q^{6k-4}e^{-2iy})(1-q^{6k-2}e^{2i(-x-y)})(1-q^{6k-4}e^{2i(x+y)}).$ (2.15)
Next, note that
$C(x,y|3\tau)=q^{2}C^{*}(x,y|3\tau)$
where $C^{*}(x,y|3\tau)$ is a function of $q^{6}$. Therefore,
$A^{3}(x,y|3\tau)-C^{3}(x,y|3\tau)=A^{3}(x,y|3\tau)-q^{6}{C^{*}}^{3}(x,y|3\tau)$
$\qquad=\prod_{j=0}^{2}\left(A(x,y|3\tau)-e^{2\pi ij/3}q^{2}C^{*}(x,y|3\tau)\right)$
$\qquad=\prod_{k=1}^{\infty}(1-q^{6k})^{6}(1-q^{18k-6}e^{6ix})(1-q^{18k-12}e^{-6ix})(1-q^{18k-6}e^{6iy})$
$\qquad\times\prod_{k=1}^{\infty}(1-q^{18k-12}e^{-6iy})(1-q^{18k-6}e^{6i(-x-y)})(1-q^{18k-12}e^{6i(x+y)}).$ (2.16)
But observe that $A^{3}(x+\pi/3,y|3\tau)-C^{3}(x+\pi/3,y|3\tau)$ has the same product representation as $A^{3}(x,y|3\tau)-C^{3}(x,y|\tau)$. Therefore,
$A^{3}(x,y|3\tau)+C^{3}(x+\pi/3,y|3\tau)=A^{3}(x+\pi/3,y|3\tau)+C^{3}(x,y|3\tau)$
or
$A^{3}(x,y|\tau)+C^{3}(x+\pi/3,y|\tau)=A^{3}(x+\pi/3,y|\tau)+C^{3}(x,y|\tau).$ (2.17)
Identity (2.17) is equivalent to Schultz’s identity (1.4). ∎
Remark 2.1.
The proof of (2.16) using (2.15) is inspired by the proof of Theorem 2.3 in [6]. The final step of proving (1.4) using the fact that the product on the right hand side of (2.16) can be written as a difference of the cube of two theta series in two ways is inspired by the proof of Corollary 4.2 in [16].
3. Analogues of (1.4)
In this section, we derive several identities analogous to (1.4). By setting $z=0$ in (2.12) and simplifying using (2.14), we deduce that
$C(x,y|\tau)-C(-x,-y|\tau)=\frac{-i}{\eta(\tau)}\vartheta_{1}(x|\tau)\vartheta_{1}(y|\tau)\vartheta_{1}(x+y|\tau),$ (3.1)
since $\vartheta_{1}(0|3\tau)=0$. Next, set $z=\pi/3$ in (2.12). Note that $\vartheta_{1}(\pi|3\tau)=0$ and therefore, simplifying using (2.14), we deduce that
$C(x,y|\tau)-\omega^{2}C(-x,-y|\tau)=\frac{-i\omega}{\eta(\tau)}\vartheta_{1}\left(\frac{\pi}{3}+x|\tau\right)\vartheta_{1}\left(\frac{\pi}{3}+y|\tau\right)\vartheta_{1}\left(\frac{\pi}{3}-x-y|\tau\right).$ (3.2)
where $\omega=\exp(2\pi i/3).$ From (3.1) and (3.2), we find that
$C(x,y|\tau)$ $=\frac{-\omega}{\sqrt{3}\eta(\tau)}\vartheta_{1}(x|\tau)\vartheta_{1}(y|\tau)\vartheta_{1}(x+y|\tau)$ (3.3)
$\quad+\frac{1}{\sqrt{3}\eta(\tau)}\vartheta_{1}\left(\frac{\pi}{3}+x|\tau\right)\vartheta_{1}\left(\frac{\pi}{3}+y|\tau\right)\vartheta_{1}\left(\frac{\pi}{3}-x-y|\tau\right).$
Replacing $(x,y)$ by (x,y)(-x,-y) in (3.2) and simplifying, we find that
$C(x,y|\tau)-\omega C(-x,-y|\tau)=\frac{i\omega^{2}}{\eta(\tau)}\vartheta_{1}\left(\frac{\pi}{3}-x|\tau\right)\vartheta_{1}\left(\frac{\pi}{3}-y|\tau\right)\vartheta_{1}\left(\frac{\pi}{3}+x+y|\tau\right).$ (3.4)
Using (2.2), we deduce that
$\vartheta_{1}(z|\tau)\vartheta_{1}\left(\frac{\pi}{3}+z|\tau\right)\vartheta_{1}\left(\frac{\pi}{3}-z|\tau\right)=\frac{\eta^{3}(\tau)}{\eta(3\tau)}\vartheta_{1}(3z|3\tau).$ (3.5)
Multiplying (3.1), (3.2) and (3.4) and using (3.5), we immediately find that
$C^{3}(x,y|\tau)-C^{3}(-x,-y|\tau)=-i\frac{\eta^{6}(\tau)}{\eta^{3}(3\tau)}\vartheta_{1}(3x|3\tau)\vartheta_{1}(3y|3\tau)\vartheta_{1}(3x+3y|3\tau).$ (3.6)
The right-hand side of (3.6) is invariant under the change $x\to x+\pi/3$ and therefore, the following new analogue of (2.17) holds:
Theorem 3.1.
Let $C(x,y|\tau)$ be given by (2.11). Then
$C^{3}(x,y|\tau)-C^{3}(-x,-y|\tau)=C^{3}(x+\pi/3,y|\tau)-C^{3}(-x-\pi/3,-y|\tau).$ (3.7)
Let $x=y=\pi\tau/3$ in (3.6). Using (2.14), we find that
$q^{2}C^{3}(-\pi\tau/3,-\pi\tau/3|\tau)-q^{2}C^{3}(\pi\tau/3,\pi\tau/3|\tau)=\frac{\eta^{9}(\tau)}{\eta^{3}(3\tau)}.$ (3.8)
Using the definition $C(x,y|\tau)$ given by (2.11) and simplifying, we deduce that (3.8) is equivalent to
$\left(\sum_{m,n=-\infty}^{\infty}q^{2(m^{2}+mn+n^{2})}\right)^{3}-\left(\sum_{m,n=-\infty}^{\infty}q^{2(m^{2}+mn+n^{2}+m+n+1/3)}\right)^{3}=\frac{\eta^{9}(\tau)}{\eta^{3}(3\tau)}.$ (3.9)
Replacing $z$ by $z+(\pi+\pi\tau)/2$ in (3.5), we find that
$\vartheta_{3}(z|\tau)\vartheta_{3}\left(\frac{\pi}{3}+z|\tau\right)\vartheta_{3}\left(\frac{\pi}{3}-z|\tau\right)=\frac{\eta^{3}(\tau)}{\eta(3\tau)}\vartheta_{3}(3z|3\tau).$ (3.10)
Equating the constant terms on both sides, we find that
$\frac{\eta^{3}(\tau)}{\eta(3\tau)}=\sum_{m,n=-\infty}^{\infty}\omega^{m-n}q^{2(m^{2}+mn+n^{2})}.$ (3.11)
Combining (3.11) and (3.9), we arrive at (1.1). This also shows that (3.7) may be viewed as a new generalization of (1.1).
Remark 3.1.
For an elegant proof of (3.11), see [6, Proposition 2.2(i)].
Setting $x=y$ in (3.1), we find that
$\sum_{m,n=-\infty}^{\infty}q^{2(m^{2}+mn+n^{2}+m+n+1/3)}\sin(2(3m+3n+2)x)=\frac{-1}{2\eta(\tau)}\vartheta_{1}^{2}(x|\tau)\vartheta_{1}(2x|\tau).$ (3.12)
Let $x=\pi/3$ in (3.12) and using
$\vartheta_{1}(\pi/3|\tau)=\sqrt{3}\eta(3\tau),$
we deduce that
$\sum_{m,n=-\infty}^{\infty}q^{2(m^{2}+mn+n^{2}+m+n+1/3)}=3\frac{\eta^{3}(3\tau)}{\eta(\tau)}.$
Setting $x=y=\pi/9$ in (3.6) and again using
$\vartheta_{1}(\pi/3|\tau)=\sqrt{3}\eta(3\tau),$
we deduce that
$\omega^{2}\left(\sum_{m,n=-\infty}^{\infty}q^{2(m^{2}+mn+n^{2}+m+n)}\omega^{m+n}\right)^{3}-\omega^{-2}\left(\sum_{m,n=-\infty}^{\infty}q^{2(m^{2}+mn+n^{2}+m+n)}\omega^{-m-n}\right)^{3}$
$=-3\sqrt{3}i\prod_{n=1}^{\infty}\frac{(1-q^{2n})^{6}(1-q^{18n})^{3}}{(1-q^{6n})^{3}}.$ (3.13)
Identity (3.13) is of the form
$aX^{3}+bY^{3}=cZ^{3}$ (3.14)
where $X,Y,Z$ are weight one modular forms and $a,b,c\in\mathbf{C}$. Note that Borweins’ identity (1.1) also satisfies (3.14) with $a=b=c=1$.
In the remaining of this section, we continue our search for identities similar to (1.4). Recall from [18, (5), p. 205] that if $M$ is a real symmetric $f\times f$ matrix, $\mathbf{m}\in\mathbf{Z}^{f}$ written as a column vector, $\mathbf{x}$ is a column vector with $f$ complex variable components $x_{1},x_{2},\cdots,x_{f}$, then for $\text{Im}\,\tau>0$, we have
$\sum_{\mathbf{m}\in\mathbf{Z}^{f}}e^{\pi i\tau(\mathbf{m}+\mathbf{x})^{t}M(\mathbf{m}+\mathbf{x})}=\dfrac{1}{(\sqrt{-i\tau})^{f}|\text{det}(M)|^{1/2}}\sum_{\mathbf{m}\in\mathbf{Z}^{f}}e^{-\pi i(\mathbf{m}^{t}M^{-1}\mathbf{m})/\tau+2\pi i\mathbf{m}^{t}\mathbf{x}}.$ (3.15)
Let $f=2$, $M=\begin{pmatrix}2&1\\ 1&2\end{pmatrix}$, $\mathbf{m}=(m,n)^{t}$ and $\mathbf{x}=(x/(\pi\tau),y/(\pi\tau))^{t}$ in (3.15). We deduce immediately that if
$A^{\dagger}(x,y|\tau)=\sum_{m,n\in\mathbf{Z}}q^{2(m^{2}-mn+n^{2})}e^{2imx}e^{2iny},$ (3.16)
then
$A(x,y|\tau)=\dfrac{1}{-i\tau\sqrt{3}}\exp\left(-\dfrac{2i(x^{2}+xy+y^{2})}{\pi\tau}\right)A^{\dagger}\left(\dfrac{x}{\tau},\dfrac{y}{\tau}\bigg|-\dfrac{1}{3\tau}\right),$
which is equivalent to
$A\left(\dfrac{x}{3\tau},\dfrac{y}{3\tau}\bigg|-\dfrac{1}{3\tau}\right)=-i\sqrt{3}\tau\exp\left(\dfrac{2i(x^{2}+xy+y^{2})}{3\pi\tau}\right)A^{\dagger}\left(-x,-y|\tau\right).$ (3.17)
Observe that if we replace $m$ by $m+n$ in (3.16), then
$A^{\dagger}(x,y|\tau)=\sum_{m,n\in\mathbf{Z}}q^{2(m^{2}+mn+n^{2})}e^{2i(m+n)x}e^{2iny}.$ (3.18)
From now on, we will use (3.18) instead of (3.16). Similarly, if
$C^{\dagger}(x,y|\tau)=\sum_{m,n\in\mathbf{Z}}\omega^{m+n}q^{2(m^{2}-mn+n^{2})}e^{2imx}e^{2iny},$ (3.19)
then
$C\left(\dfrac{x}{3\tau},\dfrac{y}{3\tau}\bigg|-\dfrac{1}{3\tau}\right)=-i\sqrt{3}\tau\exp\left(\dfrac{2i(x^{2}+xy+y^{2})}{3\pi\tau}\right)C^{\dagger}\left(-x,-y|\tau\right).$ (3.20)
Observe that if we replace $m$ by $m+n$ in (3.16), then
$C^{\dagger}(x,y|\tau)=\sum_{m,n\in\mathbf{Z}}\omega^{m-n}q^{2(m^{2}+mn+n^{2})}e^{2i(m+n)x}e^{2iny}.$
Using (3.17), (3.20) and the transformation formula [22, p. 475], [21, p. 35]
$\vartheta_{j}\left(\frac{z}{\tau}\Big|-\frac{1}{\tau}\right)=\sqrt{-i\tau}\exp\left(\frac{iz^{2}}{\pi\tau}\right)\vartheta_{j}(z|\tau),$ (3.21)
for $j=1,3$, we deduce the following identities from Theorem 2.1.
Theorem 3.2.
The following identities hold:
$3\vartheta_{1}(z+x|\tau)\vartheta_{1}(z+y|\tau)\vartheta_{1}(z-x-y|\tau)=A^{\dagger}\left(-x,-y\Big|\frac{\tau}{3}\right)\vartheta_{1}\left(z\Big|\frac{\tau}{3}\right)$
$\qquad\qquad-C^{\dagger}\left(x,y\Big|\frac{\tau}{3}\right)\vartheta_{1}\left(z-\frac{\pi}{3}\Big|\frac{\tau}{3}\right)-C^{\dagger}\left(-x,-y\Big|\frac{\tau}{3}\right)\vartheta_{1}\left(z+\frac{\pi}{3}\Big|\frac{\tau}{3}\right),$
$3\vartheta_{3}(z+x|\tau)\vartheta_{3}(z+y|\tau)\vartheta_{3}(z-x-y|\tau)=A^{\dagger}\left(-x,-y\Big|\frac{\tau}{3}\right)\vartheta_{3}\left(z\Big|\frac{\tau}{3}\right)$
$\qquad\qquad+C^{\dagger}\left(x,y\Big|\frac{\tau}{3}\right)\vartheta_{3}\left(z-\frac{\pi}{3}\Big|\frac{\tau}{3}\right)+C^{\dagger}\left(-x,-y\Big|\frac{\tau}{3}\right)\vartheta_{3}\left(z+\frac{\pi}{3}\Big|\frac{\tau}{3}\right),$
$\left(C^{\dagger}(x,y|\tau)\right)^{3}-\left(C^{\dagger}(-x,-y|\tau)\right)^{3}$
$\qquad\qquad=q^{2}e^{4ix+2iy}\left(C^{\dagger}(x+\pi\tau,y|\tau)\right)^{3}-q^{2}e^{4ix+2iy}\left(C^{\dagger}(-x-\pi\tau,-y|\tau)\right)^{3}$
and
$\left(A^{\dagger}(x,y|\tau)\right)^{3}-\left(C^{\dagger}(x,y|\tau)\right)^{3}$
$\qquad\qquad=q^{2}e^{4ix+2iy}\left(A^{\dagger}(x+\pi\tau,y|\tau)\right)^{3}-q^{2}e^{4ix+2iy}\left(C^{\dagger}(x+\pi\tau,y|\tau)\right)^{3}.$
4. Second proof of (1.4)
We will now give a second proof of (2.15), which leads to a second proof of (1.4).
Second proof of (2.15).
First, we observe that
$\{(m,n)|m,n\in\mathbf{Z}\}=\{(k+\ell,k-\ell)|k,\ell\in\mathbf{Z}\}\cup\{(k+\ell+1,k-\ell)|k,\ell\in\mathbf{Z}\}.$
Let $\gamma=1/2+\sqrt{-3}/2$. Note that $\gamma$ is a sixth-root of unity and $\gamma^{2}=\omega$. Let
$N(m+n\gamma)=(m+n\gamma)(m+n\overline{\gamma})=m^{2}+mn+n^{2}.$
Therefore,
$\sum_{m,n\in\mathbf{Z}}$ $q^{m^{2}+mn+n^{2}}e^{iu(m+n)}e^{iv(m-n)}$
$=\sum_{m,n\in\mathbf{Z}}q^{N(m+n\gamma)}e^{iu(m+n)}e^{iv(m-n)}$
$=\sum_{k,\ell\in\mathbf{Z}}q^{N((k+\ell)+(k-\ell)\gamma)}e^{2kiu}e^{2\ell iv}+\sum_{k,\ell\in\mathbf{Z}}q^{N((k+\ell+1)+(k-\ell)\gamma)}e^{iu(2k+1)}e^{iv(2\ell+1)}$
$=\sum_{k,\ell\in\mathbf{Z}}q^{3k^{2}+\ell^{2}}e^{2iku}e^{2i\ell v}+\sum_{k,\ell\in\mathbf{Z}}q^{3(k+1/2)^{2}+(\ell+1/2)^{2}}e^{iu(2k+1)}e^{iv(2\ell+1)}$
$=\vartheta_{3}(u|3\tau)\vartheta_{3}(v|\tau)+\vartheta_{2}(u|3\tau)\vartheta_{2}(v|\tau)$
or
$\sum_{m,n\in\mathbf{Z}}q^{2(m^{2}+mn+n^{2})}e^{2iu(m+n)}e^{2iv(m-n)}=\vartheta_{3}(2u|6\tau)\vartheta_{3}(2v|2\tau)+\vartheta_{2}(2u|6\tau)\vartheta_{2}(2v|2\tau).$ (4.1)
Note that by the transformation $u=3(x+y)/2$ and v=(xy)/2v=(x-y)/2 in (4.1) and using (2.10), we deduce that
$A(x,y|\tau)=\vartheta_{3}(3x+3y|6\tau)\vartheta_{3}(x-y|2\tau)+\vartheta_{2}(3x+3y|6\tau)\vartheta_{2}(x-y|2\tau).$ (4.2)
Identity (4.2) shows that in order to prove (2.15), it suffices to show that
$\vartheta_{3}(6u|6\tau)\vartheta_{3}(2v|2\tau)+\vartheta_{2}(6u|6\tau)\vartheta_{2}(2v|2\tau)$
$\qquad-q^{2/3}e^{4iu}\left(\vartheta_{3}(6u+2\pi\tau|6\tau)\vartheta_{3}(2v|2\tau)+\vartheta_{2}(6u+2\pi\tau|6\tau)\vartheta_{2}(2v|2\tau)\right)$
$=-e^{4iu}q^{2/3}\dfrac{i}{\eta(\tau)}\vartheta_{1}(2u-2\pi\tau/3|\tau)\vartheta_{1}(u+v-\pi\tau/3|\tau)\vartheta_{1}(u-v-\pi\tau/3|\tau).$ (4.3)
We will now show that (4.3) follows from an identity found by H.H. Chan, S. Cooper and P.C. Toh [9, Theorem 3.1] given by
$G_{2}(2x|2\tau)\vartheta_{2}(2y|2\tau)+G_{3}(2x|2\tau)\vartheta_{3}(2y|2\tau)=\dfrac{1}{\eta(\tau)}\vartheta_{1}(2x|\tau)\vartheta_{1}(x+y|\tau)\vartheta_{1}(x-y|\tau),$ (4.4)
where
$G_{2}(2z|2\tau)=2\sum_{\alpha\equiv 1\!\!\!\!\!\!\pmod{6}}\!\!\!\!\!\!q^{\alpha^{2}/6}\sin{2\alpha z}$
and
$G_{3}(2z|2\tau)=2\sum_{\alpha\equiv-2\!\!\!\!\!\!\pmod{6}}\!\!\!\!\!\!q^{\alpha^{2}/6}\sin{2\alpha z}.$
Using $\sin{u}=\dfrac{e^{iu}-e^{-iu}}{2i}$, we find that
$G_{2}(2u|2\tau)=\dfrac{1}{i}\left(G_{2,1}(2u|2\tau)-G_{2,2}(2u|2\tau)\right)$
where
$G_{2,1}(2u|2\tau)=\sum_{k=-\infty}^{\infty}q^{6k^{2}+2k+1/6}e^{2iu(6k+1)}$
and
$G_{2,2}(2u|2\tau)=\sum_{k=-\infty}^{\infty}q^{6k^{2}+2k+1/6}e^{-2iu(6k+1)}.$
Next, replace $u$ by $u-\pi\tau/3$. One can verify that
$G_{2,1}(2u-2\pi\tau/3|2\tau)=e^{8iu}\vartheta_{2}(6u+2\pi\tau|6\tau)$
and
$G_{2,2}(2u-2\pi\tau/3|2\tau)=q^{-2/3}e^{4iu}\vartheta_{2}(6u|6\tau).$
Similarly, if we write
$G_{3}(2u|2\tau)=\dfrac{1}{i}\left(G_{3,1}(2u|2\tau)-G_{3,2}(2u|2\tau)\right)$
where
$G_{3,1}(2u|2\tau)=\sum_{k=-\infty}^{\infty}q^{6k^{2}-4k+2/3}e^{2iu(6k-2)}$
and
$G_{3,2}(2u|2\tau)=\sum_{k=-\infty}^{\infty}q^{6k^{2}-4k+2/3}e^{-2iu(6k-2)},$
we find that
$G_{3,1}(2u-2\pi\tau/3|2\tau)=e^{8iu}\vartheta_{3}(6u+2\pi\tau|6\tau)$
and
$G_{3,2}(2u-2\pi\tau/3|2\tau)=q^{-2/3}e^{4iu}\vartheta_{3}(6u|6\tau).$
Substituting the above into (4.4), we arrive at (4.3). ∎
5. Proof of (1.6) and its analogues
Recently, Chan, Chan and Sole discover the identities [8, (1.5),(1.6)],
$\left(\sum_{m,n=-\infty}^{\infty}(-1)^{m+n}q^{dm^{2}+bmn+dn^{2}}\right)^{2}+\left(2\sum_{m,n=-\infty}^{\infty}q^{2(d(m+1/2)^{2}+b(m+1/2)n+dn^{2})}\right)^{2}$
$\qquad\qquad\qquad\qquad=\left(\sum_{m,n=-\infty}^{\infty}q^{dm^{2}+bmn+dn^{2}}\right)^{2}$ (5.1)
and
$\left(\sum_{m,n=-\infty}^{\infty}(-1)^{m+n}q^{bm^{2}+bmn+dn^{2}}\right)^{2}+\left(\sum_{m,n=-\infty}^{\infty}q^{2(b(m+1/2)^{2}+b(m+1/2)n+dn^{2})}\right)^{2}$
$\qquad\qquad\qquad\qquad=\left(\sum_{m,n=-\infty}^{\infty}q^{2(bm^{2}+bmn+dn^{2})}\right)^{2},$ (5.2)
where $b$ and $d$ satisfy certain conditions. It is natural to ask if there are any two-variable extensions for (5.1) and (5.2) similar to the generalization for (1.2) and (1.1). The answer is affirmative.
The following theorems provide generalizations of (5.1) and (5.2), respectively.
Theorem 5.1.
Suppose $b$ and $d$ are integers satisfying 2d<b<2d-2d<b<2d. Then
$\left(\sum_{m,n\in\mathbf{Z}}(-1)^{n-m}q^{d(m+1/2)^{2}+b(m+1/2)(n+1/2)+d(n+1/2)^{2}}x^{m+1/2}y^{n+1/2}\right)^{2}$
$+\left(\sum_{m,n\in\mathbf{Z}}q^{dm^{2}+bmn+dn^{2}}x^{m}y^{n}\right)^{2}=\left(\sum_{m,n\in\mathbf{Z}}(-1)^{n-m}q^{dm^{2}+bmn+dn^{2}}x^{m}y^{n}\right)^{2}$
$+\left(\sum_{m,n\in\mathbf{Z}}q^{d(m+1/2)^{2}+b(m+1/2)(n+1/2)+d(n+1/2)^{2}}x^{m+1/2}y^{n+1/2}\right)^{2}.$ (5.3)
Theorem 5.2.
Suppose $\alpha$ and $\beta$ are integers satisfying $0<\alpha<4\beta$. Then
$\left(\sum_{m,n\in\mathbf{Z}}(-1)^{n}q^{\alpha(m+1/2)^{2}+\alpha(m+1/2)n+\beta n^{2}}x^{m+1/2}y^{n}\right)^{2}$
$+\left(\sum_{m,n\in\mathbf{Z}}q^{\alpha m^{2}+\alpha mn+\beta n^{2}}x^{m}y^{n}\right)^{2}=\left(\sum_{m,n\in\mathbf{Z}}(-1)^{n}q^{\alpha m^{2}+\alpha mn+\beta n^{2}}x^{m}y^{n}\right)^{2}$
$\qquad+\left(\sum_{m,n\in\mathbf{Z}}q^{\alpha(m+1/2)^{2}+\alpha(m+1/2)n+\beta n^{2}}x^{m+1/2}y^{n}\right)^{2}.$ (5.4)
Note that (1.6) follows from (5.3) when $d=1$ and $b=0$. We derive Theorem 5.1 from the following result.
Theorem 5.3.
Let
$A_{d,b}(u,v|\tau)=\sum_{m,n\in\mathbf{Z}}q^{dm^{2}+bmn+dn^{2}}e^{2iu(m+n)}e^{2iv(m-n)}$ (5.5)
and
$C_{d,b}(u,v|\tau)$
$=\sum_{m,n\in\mathbf{Z}}q^{dm^{2}+bmn+dn^{2}+(d+b/2)m+(d+b/2)n+d/2+b/4}e^{2iu(m+n+1)}e^{2iv(m-n)}.$ (5.6)
Let $b$ and $d$ be integers satisfying 2d<b<2d-2d<b<2d so that both double sums converge. Then
$A_{d,b}(u,v|4\tau)-C_{d,b}(u,v|4\tau)=\vartheta_{4}(u|(2d+b)\tau)\vartheta_{4}(v|(2d-b)\tau).$ (5.7)
Moreover,
$A_{d,b}^{2}(u,v|\tau)-C_{d,b}^{2}(u,v|\tau)=A_{d,b}^{2}(u,v+\pi/2|\tau)-C_{d,b}^{2}(u,v+\pi/2|\tau).$ (5.8)
Similarly, we obtain Theorem 5.2 from the following.
Theorem 5.4.
Let
$\widehat{A}_{\alpha,\beta}(u,v|\tau)$ $=\sum_{m,n\in\mathbf{Z}}q^{\alpha m^{2}+\alpha mn+\beta n^{2}}e^{2iu(2m+n)}e^{2ivn}$ (5.9)
and
$\widehat{C}_{\alpha,\beta}(u,v|\tau)$ $=\sum_{m,n\in\mathbf{Z}}q^{\alpha m^{2}+\alpha mn+\beta n^{2}+\alpha m+\alpha n/2+\alpha/4}e^{2iu(2m+n+1)}e^{2ivn}.$ (5.10)
Let $0<\alpha<4\beta$ so that both double sums converge. Then
$\widehat{A}_{\alpha,\beta}(u,v|4\tau)-\widehat{C}_{\alpha,\beta}(u,v|4\tau)=\vartheta_{4}(u|\alpha\tau)\vartheta_{4}(v|(4\beta-\alpha)\tau).$ (5.11)
Moreover,
$\widehat{A}_{\alpha,\beta}^{2}(u,v|\tau)-\widehat{C}_{\alpha,\beta}^{2}(u,v|\tau)=\widehat{A}_{\alpha,\beta}^{2}(u,v+\pi/2|\tau)-\widehat{C}_{\alpha,\beta}^{2}(u,v+\pi/2|\tau).$ (5.12)
Proof of Theorem 5.3.
Every integer pair $(m,n)\in\mathbf{Z}^{2}$ corresponds uniquely to a pair $(k,\ell)\in\mathbf{Z}^{2}$ satisfying $k\equiv\ell\pmod{2}$ via the transformation
$m=\frac{k+\ell}{2}\quad\text{and}\quad n=\frac{k-\ell}{2}.$
Therefore, we may express a double sum over $(m,n)\in\mathbf{Z}^{2}$ as
$\sum_{m,n\in\mathbf{Z}}F(m,n)$ $=\sum_{\begin{subarray}{c}k,\ell\in\mathbf{Z}\\ k\equiv\ell\pmod{2}\end{subarray}}F\left(\frac{k+\ell}{2},\frac{k-\ell}{2}\right)$
$=\sum_{k,\ell\in\mathbf{Z}}F\left(k+\ell,k-\ell\right)+\sum_{k,\ell\in\mathbf{Z}}F\left(k+\ell+1,k-\ell\right),$
by considering $k,\ell$ both even and $k,\ell$ both odd.
Applying this change of summation indices on $A_{d,b}^{2}(u,v|\tau)$ and $C_{d,b}^{2}(u,v|\tau)$, respectively, we find that
$A_{d,b}(u,v|\tau)$ $=\vartheta_{3}(2u|(2d+b)\tau)\vartheta_{3}(2v|(2d-b)\tau)$
$\quad+\vartheta_{2}(2u|(2d+b)\tau)\vartheta_{2}(2v|(2d-b)\tau)$ (5.13)
and
$C_{d,b}(u,v|\tau)$ $=\vartheta_{2}(2u|(2d+b)\tau)\vartheta_{3}(2v|(2d-b)\tau)$
$\quad+\vartheta_{3}(2u|(2d+b)\tau)\vartheta_{2}(2v|(2d-b)\tau).$ (5.14)
Taking the difference between (5.13) and (5.14), we find that
$A_{d,b}(u,v|4\tau)-C_{d,b}(u,v|4\tau)$ $=\left[\vartheta_{3}(2u|4(2d+b)\tau)-\vartheta_{2}(2u|4(2d+b)\tau)\right]$
$\quad\times\left[\vartheta_{3}(2v|4(2d-b)\tau)-\vartheta_{2}(2v|4(2d-b)\tau)\right]$
$=\vartheta_{4}(u|(2d+b)\tau)\vartheta_{4}(v|(2d-b)\tau),$
since
$\vartheta_{3}(2z|4\tau)-\vartheta_{2}(2z|4\tau)=\vartheta_{4}(z|\tau),$
which follows directly from the definitions of the theta functions. This proves (5.7).
Next, we prove (5.8). It is clear from the definitions of $A_{d,b}(u,v)$ and $C_{d,b}(u,v)$, (5.5) and (5.6), that
$A_{d,b}(u+\pi/2,v+\pi/2|\tau)$ $=A_{d,b}(u,v|\tau)$ (5.15)
and
$C_{d,b}(u+\pi/2,v+\pi/2|\tau)$ $=-C_{d,b}(u,v|\tau).$ (5.16)
Replacing $u$ and $v$ by $u+\pi/2$ and $v+\pi/2$ (5.7) respectively, followed by using (5.15) and (5.16), we obtain
$A_{d,b}(u,v|4\tau)+C_{d,b}(u,v|4\tau)=\vartheta_{4}(u+\pi/2|(2d+b)\tau)\vartheta_{4}(v+\pi/2|(2d-b)\tau).$
Multiplying this equation by (5.7), we then obtain
$A_{d,b}^{2}(u,v|4\tau)-C_{d,b}^{2}(u,v|4\tau)$ $=\vartheta_{4}(u|(2d+b)\tau)\vartheta_{4}(v|(2d-b)\tau)$
$\qquad\times\vartheta_{4}(u+\pi/2|(2d+b)\tau)\vartheta_{4}(v+\pi/2|(2d-b)\tau).$ (5.17)
Since the expression on the right hand side of (5.17) is invariant when $v$ is replaced by $v+\pi/2$, we obtain
$A_{d,b}^{2}(u,v|4\tau)-C_{d,b}^{2}(u,v|4\tau)=A_{d,b}^{2}(u,v+\pi/2|4\tau)-C_{d,b}^{2}(u,v+\pi/2|4\tau)$
and the proof of Theorem 5.3 is complete.
Proof of Theorem 5.1.
Set $x=e^{2i(u+v)}$ and $y=e^{2i(u-v)}$ in (5.3). With these substitutions, (5.3) becomes equivalent to (5.8). Since (5.8) holds, the theorem follows immediately. ∎
Next, we show that (5.8) is a two-variable generalization of (5.1).
From (5.14), we find that
$C_{d,b}(0,0|\tau)-C_{d,b}(0,\pi/2|\tau)=2\vartheta_{3}(0|(2d+b)\tau)\vartheta_{2}(0|(2d-b)\tau)$ (5.18)
and
$C_{d,b}(0,0|\tau)+C_{d,b}(0,\pi/2|\tau)=2\vartheta_{2}(0|(2d+b)\tau)\vartheta_{3}(0|(2d-b)\tau).$ (5.19)
Multiplying (5.18) by (5.19), we arrive at
$C_{d,b}^{2}(0,0|\tau)-C_{d,b}^{2}(0,\pi/2|\tau)$
$=4\vartheta_{2}(0|(2d+b)\tau)\vartheta_{3}(0|(2d+b)\tau)\vartheta_{2}(0|(2d-b)\tau)\vartheta_{3}(0|(2d-b)\tau)$
$=\vartheta_{2}^{2}(0|(d+b/2)\vartheta_{2}^{2}(0|(d-b/2)\tau),$
where we have used [7, p.58]
$2\vartheta_{2}(0|2\tau)\vartheta_{3}(0|2\tau)=\vartheta_{2}^{2}(0|\tau).$ (5.20)
Applying the change of summation indices $m=(k+\ell)/2$ and $n=(k-\ell)/2$, as in the proof of Theorem 5.3, we find that
$2\sum_{m,n\in\mathbf{Z}}q^{2(d(m+1/2)^{2}+b(m+1/2)n+dn^{2})}$
$=2\sum_{k,\ell\in\mathbf{Z}}q^{2(2d+b)(k+1/4)^{2}+(2d-b)(\ell+1/4)^{2}}+2\sum_{k,\ell\in\mathbf{Z}}q^{2(2d+b)(k+3/4)^{2}+(2d-b)(\ell+3/4)^{2}}$
$=4\sum_{k,\ell\in\mathbf{Z}}q^{2(2d+b)(k+1/4)^{2}+(2d-b)(\ell+1/4)^{2}}$
$=\vartheta_{2}(0|(d+b/2)\vartheta_{2}(0|(d-b/2)\tau),$
where the last equality follows from the following computations:
$2\sum_{k\in\mathbf{Z}}q^{4(k+1/4)^{2}}=2\sum_{k\in\mathbf{Z}}q^{4(k-1/4)^{2}}=2\sum_{k\in\mathbf{Z}}q^{4(k+3/4)^{2}}$
$=\sum_{k\in\mathbf{Z}}q^{(4k+1)^{2}/4}+\sum_{k\in\mathbf{Z}}q^{(4k+3)^{2}/4}=\sum_{k\in\mathbf{Z}}q^{(2k+1)^{2}/4}=\vartheta_{2}(0|\tau).$
Therefore,
$C_{d,b}^{2}(0,0|\tau)-C_{d,b}^{2}(0,\pi/2|\tau)=\left(2\sum_{m,n\in\mathbf{Z}}q^{2(d(m+1/2)^{2}+b(m+1/2)n+dn^{2})}\right)^{2},$
and by setting $u=v=0$ in (5.8), we recover (5.1).
Proof of Theorem 5.4.
Instead of proving (5.11) directly, thereby replicating similar proofs to that of (5.7), we deduce (5.11) from (5.7). By replacing the summation variables $(m,n)$ by $(\ell,k-\ell)$, we find that
$\widehat{A}_{\alpha,\beta}(u,v|\tau)$ $=A_{\beta,\alpha-2\beta}(u,v|\tau)$
and
$\widehat{C}_{\alpha,\beta}(u,v|\tau)$ $=C_{\beta,\alpha-2\beta}(u,v|\tau).$
Therefore, by applying (5.7),
$\widehat{A}_{\alpha,\beta}(u,v|4\tau)-\widehat{C}_{\alpha,\beta}(u,v|4\tau)$
$=A_{\beta,\alpha-2\beta}(u,v|4\tau)-C_{\beta,\alpha-2\beta}(u,v|4\tau)$
$=\vartheta_{4}(u|\alpha\tau)\vartheta_{4}(v|(4\beta-\alpha)\tau),$
which is (5.11). Identity (5.12) also follows easily from (5.8) by nothing that
$\widehat{A}_{\alpha,\beta}^{2}(u,v|\tau)-\widehat{C}_{\alpha,\beta}^{2}(u,v|\tau)$ $=A_{\beta,\alpha-2\beta}^{2}(u,v|\tau)-C_{\beta,\alpha-2\beta}^{2}(u,v|\tau)$
$=A_{\beta,\alpha-2\beta}^{2}(u,v+\pi/2|\tau)-C_{\beta,\alpha-2\beta}^{2}(u,v+\pi/2|\tau)$
$=\widehat{A}_{\alpha,\beta}^{2}(u,v+\pi/2|\tau)-\widehat{C}_{\alpha,\beta}^{2}(u,v+\pi/2|\tau).$
Proof of Theorem 5.2.
Set $x=e^{4iu}$ and $y=e^{2i(u+v)}$ in (5.4). With these substitutions, (5.4) becomes equivalent to (5.12). Since (5.12) holds, the theorem follows immediately. ∎
Finally, we show that (5.12) is a two variable generalization of [8, (1.6)]. Comparing (5.12) and [8, (1.6)], we see that it suffices to show
$\widehat{A}_{\alpha,\beta}^{2}(0,\pi/2|2\tau)-\widehat{C}_{\alpha,\beta}^{2}(0,\pi/2|2\tau)=\left(\sum_{m,n\in\mathbf{Z}}(-1)^{m-n}q^{\alpha m^{2}+\alpha mn+\beta n^{2}}\right)^{2}.$ (5.21)
From (5.17), we see that
$\widehat{A}_{\alpha,\beta}^{2}(0,\pi/2|4\tau)-\widehat{C}_{\alpha,\beta}^{2}(0,\pi/2|4\tau)$
$=\vartheta_{4}(0|\alpha\tau)\vartheta_{4}(0|(4\beta-\alpha)\tau)\vartheta_{4}(\pi/2|\alpha\tau)\vartheta_{4}(\pi/2|(4\beta-\alpha)\tau)$
$=\vartheta_{4}(0|\alpha\tau)\vartheta_{4}(0|(4\beta-\alpha)\tau)\vartheta_{3}(0|\alpha\tau)\vartheta_{3}(0|(4\beta-\alpha)\tau)$
$=\vartheta_{4}^{2}(0|2\alpha\tau)\vartheta_{4}^{2}(0|2(4\beta-\alpha)\tau)$
by applying the identity [3, p.34]
$\vartheta_{3}(0|\tau)\vartheta_{4}(0|\tau)=\vartheta_{4}^{2}(0|2\tau).$ (5.22)
Since it is shown in [8, (3.2)] that
$\sum_{m,n\in\mathbf{Z}}(-1)^{m-n}q^{\alpha m^{2}+\alpha mn+\beta n^{2}}=\vartheta_{4}(0|\alpha\tau)\vartheta_{4}(0|(4\beta-\alpha)\tau),$
equation (5.21) then follows immediately.
Remark 5.1.
Both (5.1) and (5.2) are generalizations of (1.2). This can be seen as follows.
By setting $b=1$ and $d=0$ in (5.1), the equivalence between (5.1) and (1.2) is established once we show that
$2\sum_{m,n=-\infty}^{\infty}q^{2(m+1/2)^{2}+2n^{2}}=\sum_{m,n=-\infty}^{\infty}q^{(m+1/2)^{2}+(n+1/2)^{2}},$ (5.23)
but this is precisely identity (5.20).
Next, to show (5.2) implies (1.2), we set $b=1$ and $d=1/2$ in (5.2), and then replace the summation indices $(m,n)$ by $(k,\ell-k)$. Doing so, we arrive at
$\left(\sum_{m,n=-\infty}^{\infty}(-1)^{n}q^{(m^{2}+n^{2})/2}\right)^{2}+\left(\sum_{m,n=-\infty}^{\infty}q^{(m+1/2)^{2}+(n+1/2)^{2}}\right)^{2}$
$\qquad\qquad\qquad\qquad=\left(\sum_{m,n=-\infty}^{\infty}q^{m^{2}+n^{2}}\right)^{2}.$
Thus, it suffices to show that
$\sum_{m,n=-\infty}^{\infty}(-1)^{n}q^{m^{2}+n^{2}}=\sum_{m,n=-\infty}^{\infty}(-1)^{m+n}q^{2(m^{2}+n^{2})},$ (5.24)
which again is precisely identity (5.22).
6. Cubic analogues of (1.1)
It is natural to ask if there are cubic analogues of (5.1) and (5.2). This question leads us to the following new identities of Borweins-type.
Theorem 6.1.
Let $|q|<1$ and $\omega=e^{2\pi i/3}$. Then
$\left(\sum_{m,n\in\mathbf{Z}}\omega^{m-n}q^{m^{2}+mn+n^{2}}\right)^{3}$ $+\left(3\sum_{m,n\in\mathbf{Z}}q^{3((m+1/3)^{2}+(m+1/3)n+n^{2})}\right)^{3}$
$\qquad\qquad=\left(\sum_{m,n\in\mathbf{Z}}q^{m^{2}+mn+n^{2}}\right)^{3}$ (6.1)
and
$\left(\sum_{m,n\in\mathbf{Z}}\omega^{m-n}q^{3m^{2}+3mn+n^{2}}\right)^{3}$ $+\left(\sum_{m,n\in\mathbf{Z}}q^{3(3(m+1/3)^{2}+3(m+1/3)n+n^{2})}\right)^{3}$
$\qquad\qquad=\left(\sum_{m,n\in\mathbf{Z}}q^{3(3m^{2}+3mn+n^{2})}\right)^{3}.$ (6.2)
We note that (6.1) is the cubic analogue of (5.1) in the case $b=d=1$ and (6.2) is the cubic analogue of (5.2) in the case $b=3$ and $d=1$. These identities are different ways of expressing (1.1) but establishing their connections with (1.1) turns out to be very interesting.
We first require three identities.
Lemma 6.1.
Let $|q|<1$ and $\omega=e^{2\pi i/3}$. Then
$\sum_{m,n\in\mathbf{Z}}\omega^{m-n}q^{m^{2}+mn+n^{2}+m+n}=0$ (6.3)
$\sum_{m,n\in\mathbf{Z}}\omega^{m+n}q^{m^{2}+mn+n^{2}}=\sum_{m,n\in\mathbf{Z}}\omega^{m-n}q^{3(m^{2}+mn+n^{2})}$ (6.4)
and
$\sum_{m,n\in\mathbf{Z}}q^{(m+1/3)^{2}+(m+1/3)(n+1/3)+(n+1/3)^{2}}$
$=3\sum_{m,n\in\mathbf{Z}}q^{3((m+1/3)^{2}-(m+1/3)(n+1/3)+(n+1/3)^{2})}.$ (6.5)
Proof.
Identity (6.3) is (1.5) and the proof is presented in Section 1. Let
$N(m+n\gamma)=\left|m+n\gamma\right|^{2}=m^{2}+mn+n^{2},$
$R(m+n\gamma)=m$
and
$I(m+n\gamma)=n,$
where $\gamma=(1+\sqrt{-3})/2.$ From
$\mathcal{L}=\bigcup_{j=0}^{2}\mathcal{L}_{j}$
where
$\mathcal{L}=\{m+n\gamma|m,n\in\mathbf{Z}\}\quad\text{and}\quad\mathcal{L}_{j}=\{-2n-m+j+(2m+n)\gamma|m,n\in\mathbf{Z}\},$
we deduce that
$\sum_{\ell\in\mathcal{L}}\omega^{R(\ell)+I(\ell)}q^{N(\ell)}=\sum_{j=0}^{2}T_{j}$
where
$T_{j}=\sum_{\ell\in\mathcal{L}_{j}}\omega^{R(\ell)+I(\ell)}q^{N(\ell)}.$
Note that
$T_{0}$ $=\sum_{m,n\in\mathbf{Z}}\omega^{m-n}q^{3m^{2}+3mn+3n^{2}},$
$T_{1}$ $=\sum_{m,n\in\mathbf{Z}}\omega^{m-n+1}q^{3m^{2}+3mn+3n^{2}-3n+1}$
and
$T_{2}$ $=\sum_{m,n\in\mathbf{Z}}\omega^{m-n-1}q^{3m^{2}+3mn+3n^{2}+3n+1}.$
Replacing $n$ by $n+1$ in $T_{1}$, we deduce that
$T_{1}=\sum_{m,n\in\mathbf{Z}}\omega^{m-n}q^{3m^{2}+3mn+3n^{2}+3m+3n+1}=0,$
by (6.3). Similarly, by replacing $n$ by n1n-1 in $T_{2}$, followed by $(m,n)$ by (n,m)(-n,-m), and using (6.3), we conclude that $T_{2}=0$ and the proof of (6.4) is complete.
To prove (6.5), we observe that by direct verification that
$\sum_{m,n\in\mathbf{Z}}q^{3m^{2}+3n^{2}+3mn+3m+3n+1}=\sum_{\begin{subarray}{c}m,n\in\mathbf{Z}\\ m\equiv n+1\!\!\!\!\!\pmod{3}\end{subarray}}q^{m^{2}+mn+n^{2}}.$
The solutions to $m\equiv n+1\pmod{3}$ are $(m,n)\equiv(1,0),(-1,1)$ and $(0,-1)\pmod{3}$. Therefore,
$\sum_{\begin{subarray}{c}m,n\in\mathbf{Z}\\ m\equiv n+1\!\!\!\!\!\pmod{3}\end{subarray}}q^{m^{2}+mn+n^{2}}=C_{1,0}+C_{0,-1}+C_{-1,1},$
where
$C_{i,j}=\sum_{\begin{subarray}{c}m,n\in\mathbf{Z}\\ (m,n)\equiv(i,j)\!\!\!\!\!\pmod{3}\end{subarray}}q^{m^{2}+mn+n^{2}}.$
We will show that $C_{1,0}=C_{0,-1}=C_{-1,1}$. Note that
$C_{1,0}=C_{-1,0}=C_{0,-1}.$
To prove that $C_{-1,1}=C_{1,0}$, we observe that
$C_{-1,1}=\sum_{k,\ell\in\mathbf{Z}}q^{9k^{2}+9k\ell+9\ell^{2}-3k+3\ell+1}$ (6.6)
and
$C_{1,0}=\sum_{k,\ell\in\mathbf{Z}}q^{9k^{2}+9k\ell+9\ell^{2}+6k+3\ell+1}$
and that $C_{-1,1}=C_{1,0}$ after replacing $(k,\ell)$ by $(-k-\ell,k)$ in $C_{-1,1}.$ Therefore,
$\sum_{m,n\in\mathbf{Z}}q^{3(m+1/3)^{2}+3(m+1/3)(n+1/3)+3(n+1/3)^{2}}=3C_{-1,1}$ (6.7)
and the proof of (6.5) is complete.
Remark 6.1.
  1. 1.
    Identity (6.3) can also be proved by writing $S$ as S=ω^-1S_-1+S_0+ωS_1, where S_j = ∑_m,n∈Zm-n≡j (mod3) q^m^2+mn+n^2+m+n. It can be shown that that $S_{-1}=S_{0}=S_{1}$ so that S=(ω^-1+1+ω) S_0 = 0 since $1+\omega+\omega^{-1}=0.$ We leave the details to the reader.
  2. 2.
    Identity (6.5) is not new. The identity and its proof can be found in [25], [13, Lemma 3.13] and [11, (1.11)].
  3. 3.
    Identities (6.4) and (6.5) (in the equivalent form (6.8) below) can be viewed as cubic analogues of (5.24) and (5.23), respectively.
We are now ready to prove Theorem 6.1.
Proof of Theorem 6.1.
By replacing $(m,n)$ by $(m,m+n)$ for the series on the right hand side of (6.5), we deduce that
$\sum_{m,n\in\mathbf{Z}}q^{(m+1/3)^{2}+(m+1/3)(n+1/3)+(n+1/3)^{2}}=3\sum_{m,n\in\mathbf{Z}}q^{3((m+1/3)^{2}+(m+1/3)n+n^{2})}.$ (6.8)
Thus Borweins’ cubic identity (1.1) can be expressed as
$\left(\sum_{m,n\in\mathbf{Z}}\omega^{m-n}q^{m^{2}+mn+n^{2}}\right)^{3}$ $+\left(3\sum_{m,n\in\mathbf{Z}}q^{3((m+1/3)^{2}+(m+1/3)n+n^{2})}\right)^{3}$
$\qquad\qquad=\left(\sum_{m,n\in\mathbf{Z}}q^{m^{2}+mn+n^{2}}\right)^{3},$
which is the cubic analog of [8, eq. (1.5)] with $b=d=1$.
Replacing $(m,n)$ by $(m,m+n)$ in the second series on the left hand side and the series on the right hand side of (1.1), we find that
$\left(\sum_{m,n\in\mathbf{Z}}\omega^{m-n}q^{m^{2}+mn+n^{2}}\right)^{3}$ $+\left(\sum_{m,n\in\mathbf{Z}}q^{3(m+1/3)^{2}+3(m+1/3)n+n^{2}}\right)^{3}$
$\qquad\qquad=\left(\sum_{m,n\in\mathbf{Z}}q^{3m^{2}+3mn+n^{2}}\right)^{3},$
which, after replacing $q$ by $q^{3}$, that
$\left(\sum_{m,n\in\mathbf{Z}}\omega^{m-n}q^{3(m^{2}+mn+n^{2})}\right)^{3}$ $+\left(\sum_{m,n\in\mathbf{Z}}q^{3(3(m+1/3)^{2}+3(m+1/3)n+n^{2})}\right)^{3}$
$\qquad\qquad=\left(\sum_{m,n\in\mathbf{Z}}q^{3(3m^{2}+3mn+n^{2})}\right)^{3}.$ (6.9)
But by (6.4),
$\sum_{m,n\in\mathbf{Z}}\omega^{m-n}q^{3(m^{2}+mn+n^{2})}$ $=\sum_{m,n\in\mathbf{Z}}\omega^{m+n}q^{m^{2}+mn+n^{2}}$
$=\sum_{m,n\in\mathbf{Z}}\omega^{2m+n}q^{3m^{2}+3mn+n^{2}}$
$=\sum_{m,n\in\mathbf{Z}}\omega^{n-m}q^{3m^{2}+3mn+n^{2}},$ (6.10)
where the second equality follows by replacing $(m,n)$ by $(m,m+n)$ and the third equality follows from $\omega^{2m}=\omega^{-m}.$ Substituting (6.10) into (6.9), we arrive at (6.2). ∎
Remark 6.2.
Using (6.6), (6.7), (1.1) and the fact that
$C_{-1,1}=\sum_{m,n\in\mathbf{Z}}q^{9((m+1/3)^{2}-(m+1/3)(n+1/3)+(n+1/3)^{2})},$
we deduce that
$\left(\sum_{m,n\in\mathbf{Z}}\omega^{m-n}q^{m^{2}+mn+n^{2}}\right)^{3}$ $+\left(3\sum_{m,n\in\mathbf{Z}}q^{3((m+1/3)^{2}-(m+1/3)(n+1/3)+(n+1/3)^{2})}\right)^{3}$
$\qquad\qquad=\left(\sum_{m,n\in\mathbf{Z}}q^{m^{2}+mn+n^{2}}\right)^{3}$ (6.11)
Using (6.4) and (1.1), we arrive at
$\left(\sum_{m,n\in\mathbf{Z}}\omega^{m+n}q^{m^{2}+mn+n^{2}}\right)^{3}$ $+\left(\sum_{m,n\in\mathbf{Z}}q^{3((m+1/3)^{2}+(m+1/3)(n+1/3)+(n+1/3)^{2})}\right)^{3}$
$\qquad\qquad=\left(\sum_{m,n\in\mathbf{Z}}q^{3(m^{2}+mn+n^{2})}\right)^{3}.$ (6.12)
Identities (6.11) and (6.12) are equivalent to (1.4) and their discoveries are motivated by [8].
7. Concluding remarks
In this article, we have shown that $A(x,y|\tau)$ and $C(x,y|\tau)$ are coefficients of certain theta series identities such as (2.12). We then use (2.12) to prove (1.4). We have also connected (1.4) to identity connected to the study of Macdonald identities. Using transformation formulas from classical texts, we derive new analogues of (1.4). Motivated by (1.4) and identities discovered in [8], we found certain $X,Y,Z,W$, which are two-variable extension of series discovered in [8], that satisfy
$X^{2}+Y^{2}=Z^{2}+W^{2},$
We end the article with several new series $A,B,C$ that satisfy
$A^{3}+B^{3}=C^{3}.$
All the identities we found in that section are different ways of expressing (1.1).